OpenStudy (anonymous):

There are 7 girls in a bus. Each girl has 7 backpacks. In each backpack, there are 7 big cats. For every big cat there are 7 little cats. Question: How many legs are there in the bus?

OpenStudy (anonymous):

tricky ;)

OpenStudy (anonymous):

7x7x7

OpenStudy (anonymous):

You never stated how many legs each cat has. Lola, you're wrong.

OpenStudy (anonymous):

7 girls. 7 back packs. 7 big cats, 4 legs each. 49 little cats, 4 legs each

OpenStudy (anonymous):

frogwingspans your so clever...just answer it

OpenStudy (anonymous):

each has four so 7x7x7x4x4x4x4x4x4x4x4x4x4x4x some thing like that

OpenStudy (anonymous):

someone help me with my question for 2 medals

OpenStudy (anonymous):

humans dont have 4 legs, only two. :)

OpenStudy (anonymous):

is it 196?

OpenStudy (anonymous):

urgggggghhhhhhh.............hard to answer when your supervisor is just a few feet away...heheheh need to work now!

OpenStudy (anonymous):

I didn't calculate the numbers, but 196 seems far too little. and lmao

OpenStudy (anonymous):

oh, as we all know, there are four legs in each cat. so this must be 7x4=28 (for the little kittens) and also for the big cats then 28+28=56(for the cats) for the girls, we have 7x2=14 then add, so this is 14+56 which will give us 70 legs in all.

OpenStudy (anonymous):

wow you are smart

OpenStudy (anonymous):

Wrong. the little cats alone have 196 legs

OpenStudy (anonymous):

so its 392?

OpenStudy (anonymous):

wait, thats still wrong.

OpenStudy (anonymous):

ha ha ha naughty frog

OpenStudy (anonymous):

196?are you sure?it's not

OpenStudy (anonymous):

eugh fine I'll grind out the numbers.

OpenStudy (anonymous):

there are only 4 legs in a kitten so multiply 7 by 4.

OpenStudy (anonymous):

yeah i see.

OpenStudy (anonymous):

there are 7 kittens for every cat. so 49 kittens

OpenStudy (anonymous):

There are 7 girls in a bus. 7 * 2 = 14 Each girl has 7 backpacks. In each backpack, there are 7 big cats. = 7*4 = 28 For every big cat there are 7 little cats. = 7*7*4 = 196 total = 196 + 14 + 28 = 238

OpenStudy (anonymous):

^

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

wait... we missed something

OpenStudy (anonymous):

Each girl has 7 backpacks... so 49 backpacks Wow....

OpenStudy (anonymous):

I don't even want to touch those numbers.

OpenStudy (anonymous):

i got 1386

OpenStudy (anonymous):

that sounds right @ pat

OpenStudy (anonymous):

1386

OpenStudy (anonymous):

got the logic now

OpenStudy (anonymous):

I think we should just delete this seeing as the person who asked the question isn't even trying to learn.

OpenStudy (anonymous):

some answer goes like these 14 girl legs 1372 big cat legs 9604 little cat legs 10990 legs

OpenStudy (anonymous):

urgh...now I'm confused again :(

OpenStudy (anonymous):

let's have the 7 kittens first, so there are 28 legs for the kittens, but it is still 7 big cats so, 28*7=196 legs x 49 backpacks= 9604 legs + 14=9618 my answer is 9618

OpenStudy (anonymous):

There are 7 girls in a bus. 7*2 = 14 girls leg Each girl has 7 backpacks. 7*7 = 49 backpacks In each backpack, there are 7 big cats. 7* 49 = 63 big cats = 252 legs For every big cat there are 7 little cats. 63*7 = 441 little cats = 1764 legs total = 14 + 252 + 1764 = 2030 legs

OpenStudy (anonymous):

Just tell your teacher you're dyslexic.

OpenStudy (anonymous):

your so bad frog...

OpenStudy (anonymous):

oh yeah now I got it for sure...10990 oh yeah

OpenStudy (anonymous):

nope i'm not dyslexic, but its okay...We are free to express right?

OpenStudy (anonymous):

your ryt miss bio

OpenStudy (anonymous):

forg your not thinking

OpenStudy (anonymous):

there are 7 girls. there are 343 big cats in all. there are 2401 little cats. so 7*2=14 legs(for girls) 343*4=1372 legs(for big cats) 2401*4=9604 legs(for little cats) add everything so the answer is 10,990 legs in all, unless other legs of cat are cut.hehehe

OpenStudy (anonymous):

i think that is the answer now.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

yah you got it

OpenStudy (anonymous):

we got it:)